OpenStudy (anonymous):

line contains points (2, -2) and (-4, 1). Another line graphed in the same coordinate plane contains points (1, 2) and (-3, 4). Based on the slopes of these lines, are they parallel, perpendicular or neither?

OpenStudy (swissgirl):

Ok so first we have to the slope of both lines

OpenStudy (swissgirl):

Do you remember the formula?

OpenStudy (anonymous):

y2-y1/x-x1

OpenStudy (swissgirl):

\( (2,-2)=(x_1,y_1)\) \((-4,1)=(x_2,y_2)\)

OpenStudy (anonymous):

Parallel

OpenStudy (swissgirl):

\( \large {1-(-2) \over -4-(2)}=\frac{3}{6}=\frac{1}{2}\)

OpenStudy (swissgirl):

ughhh its -1/2 i forgot the negative

OpenStudy (anonymous):

Parellel?

OpenStudy (swissgirl):

Try to find the slope of the second line Use the same process

OpenStudy (anonymous):

I got parellel

OpenStudy (swissgirl):

Yes the lines are parallel since tehy both have the same slope but like i am assuming you have to solve it not just say its parallel